Consider the equation 3 E ≡ 5x − 2y = .
Write down equations E , respectively so that 1 E,
2 E,
3
i) E and E are inconsistent; 1
ii) E and E have a unique solution; 2
iii) E and E have infinitely many solutions.
1
Expert's answer
2019-03-28T12:09:34-0400

For a system of linear equations in two variables, exactly one of the following is true.


"\begin{matrix}
 a_1x+b_1y=c_1 \\
 a_2x+b_2y=c_2
\end{matrix}"

1. The system has no solution.


"\begin{vmatrix}
 a_1 & b_1 \\
 a_2 & b_2
\end{vmatrix}=0\ but \ \begin{vmatrix}
 c_1 & b_1 \\
 c_2 & b_2
\end{vmatrix}=\not 0\ or \begin{vmatrix}
 a_1 & c_1 \\
 a_2 & c_2
\end{vmatrix}=\not0"

"\begin{matrix}
 E : 5x-2y=3 \\
 E1: -10x+4y=5
\end{matrix}"

2. The system has exactly one solution.


"\begin{vmatrix}
 a_1 & b_1 \\
 a_2 & b_2
\end{vmatrix}=\not0"

"\begin{matrix}
 E : 5x-2y=3 \\
 E2: x+4y=5
\end{matrix}"

3. The system has infinitely many solutions.


"\begin{vmatrix}
 a_1 & b_1 \\
 a_2 & b_2
\end{vmatrix}=0 , \begin{vmatrix}
 c_1 & b_1 \\
 c_2 & b_2
\end{vmatrix}=0, \begin{vmatrix}
 a_1 & c_1 \\
 a_2 & c_2
\end{vmatrix}=0"

"\begin{matrix}
 E : 5x-2y=3 \\
 E3: -15x+6y=-9
\end{matrix}"